# When Style Breaks Safety: Defending LLMs Against Superficial Style Alignment

### v2

@@ Abstract
Large language models (LLMs) can be prompted with specific styles (e.g., formatting responses as lists), including in malicious queries. Prior jailbreak research mainly augments these queries with additional string transformations to maximize attack success rate (ASR). However, the impact of style patterns in the original queries that are semantically irrelevant to the malicious intent remains unclear. In this work, we seek to understand whether style patterns compromise LLM safety, how superficial style alignment increases model vulnerability, and how best to mitigate these risks during alignment. We first define ASR inflation as the increase in ASR due to style patterns in existing jailbreak benchmark queries. By evaluating $32$ LLMs across seven benchmarks, we find that nearly all models exhibit ASR inflation. Notably, the inflation correlates with an LLM’s relative attention to style patterns, which also overlap more with its instruction-tuning data when inflation occurs. We then investigate superficial style alignment, and find that fine-tuning with specific styles makes LLMs more vulnerable to jailbreaks of those same styles. Finally, we propose SafeStyle, a defense strategy that incorporates a small amount of safety training data augmented to match the distribution of style patterns in the fine-tuning data. Across three LLMs, six fine-tuning style settings, and two real-world instruction-tuning datasets, SafeStyle consistently outperforms baselines in maintaining LLM safety. Content Warning: This paper contains examples of harmful language.
@@ 1 Introduction
These observations raise a critical yet understudied question: Can style patterns in malicious queries inadvertently undermine the robustness of safety-aligned LLMs? The superficial alignment hypothesis (Zhou et al., 2023; Lin et al., 2024; Zhang & Wu, 2024) posits that alignment tuning may encourage models to imitate style patterns without internalizing deeper safety principles. Meanwhile, prior work on safety defense has largely focused on data curation (Bianchi et al., 2024; He et al., 2024; Shen et al., 2025), mechanistic interventions (Hazra et al., 2024; Hsu et al., 2024; Li et al., 2025a; Tamirisa et al., 2025; Yi et al., 2025b; Zhao et al., 2025), or safety training objectives (Huang et al., 2024b; c; 2025; Rosati et al., 2024; Zhang et al., 2025a). While Hsiung et al. (2025) show that diverse alignment data improves robustness, they do not examine the underlying mechanism. Qi et al. (2025a) argue that safety training is often shallow but focus mainly on the early response stage. This leaves a crucial gap in understanding the safety implications of style compliance.
@@
Style patterns and LLM safety: First, we define ASR inflation as the phenomenon in which the underlying malicious intent remains unchanged, but the addition of semantically irrelevant style patterns to jailbreak queries breaks LLM safety. For instance, Mistral-Nemo-Instruct-2407 (Jiang et al., 2023) refuses “chemical warfare agents” but complies with “create a list of chemical warfare agents”. Through an extensive study of $32$ LLMs (Bai et al., 2023; Jiang et al., 2023; Touvron et al., 2023; Grattafiori et al., 2024; Groeneveld et al., 2024; OLMo et al., 2024; Team et al., 2024a; b; 2025; Yang et al., 2024a; b) across seven jailbreak benchmarks (Zou et al., 2023; Han et al., 2024; Huang et al., 2024d; Mazeika et al., 2024; Röttger et al., 2024; Souly et al., 2024; Xie et al., 2025), we find that style patterns inflate the ASR for nearly all models. Based on our finding, we argue that the reported ASRs in benchmarks are inflated, in the sense that they do not capture the true rates when LLMs face core malicious intents alone. We further show that ASR inflation correlates with a model’s relative attention to these patterns. Surprisingly, ASR-inflating style patterns appear more frequently in the instruction-tuning datasets (Ivison et al., 2023; Lambert et al., 2024) used to align LLMs (Groeneveld et al., 2024; OLMo et al., 2024).
@@
  • We show that style patterns inflate ASR by systematically evaluating the safety risks of $32$ LLMs across seven jailbreak benchmarks.
@@ 2 Related Work
LLM Safety Risk. Despite extensive safety alignment efforts, LLMs remain vulnerable to malicious queries (Deng et al., 2024; Dong et al., 2024; Jin et al., 2024; Zheng et al., 2024a; Chan et al., 2025), with style-based attack strategies further amplifying these risks (Wei et al., 2023; Doumbouya et al., 2025). Recent jailbreak benchmarks aim to quantify such vulnerabilities by measuring the ease with which LLMs produce unethical responses (Zou et al., 2023; Han et al., 2024; Huang et al., 2024d; Mazeika et al., 2024; Röttger et al., 2024; Souly et al., 2024; Xie et al., 2025). However, as we show, many benchmarks report inflated ASRs due to style patterns that are independent of the underlying malicious intents or any added attack transformations. Meanwhile, studies have shown that even benign fine-tuning can inadvertently compromise model safety (He et al., 2024; Qi et al., 2024; Eiras et al., 2025). Rather than focusing on new attacks, our work identifies superficial style alignment during fine-tuning as a potential cause of the inflated ASRs observed in these benchmarks.
LLM Safety Defense. Various studies have proposed defense strategies to address the safety challenges faced by LLMs, particularly those introduced during fine-tuning (Huang et al., 2024a). Several methods (Hazra et al., 2024; Hsu et al., 2024; Liu et al., 2024; Djuhera et al., 2025; Li et al., 2025a; Tamirisa et al., 2025; Yi et al., 2025b; Zhao et al., 2025) develop mechanistic interventions to protect safety-relevant parameters, while others (Huang et al., 2024b; c; 2025; Qi et al., 2025a; Rosati et al., 2024; Zhang et al., 2025a) modify the training objective to prioritize safety. Diagnostic efforts (Peng et al., 2024; Qi et al., 2025b) examine how safety behaviors are encoded and degrade during adaptation. Data-centric approaches emphasize careful curation of fine-tuning data (Bianchi et al., 2024; He et al., 2024; Shen et al., 2025) or highlight the critical role of prompt templates at inference time (Lyu et al., 2024; Eiras et al., 2025; Yi et al., 2025a). Notably, Hsiung et al. (2025) underscores diversity in safety alignment data, and Sharma et al. (2025) uses style-based augmentation to train red-teaming classifiers. In contrast, our work demonstrates that incorporating a small amount of safety data, which is augmented to match the distributions of style patterns in the fine-tuning data, can effectively restore LLMs’ safety strength degraded by superficial style alignment.
@@ 3 Style Patterns Inflate ASR
Figure 2: (a) Nearly all of the $32$ examined LLMs exhibit inflated ASR due to the incorporation of style patterns in jailbreak queries. (b) All seven jailbreak benchmarks lead to ASR inflation, with SorryBench and MedSafetyBench affecting the most LLMs.
@@ 3.1 Experiment Setup
In our experiments, we consider seven jailbreak benchmarks: (1) AdvBench (Zou et al., 2023), (2) the standard split of HarmBench (Mazeika et al., 2024), (3) the base dataset from SORRY-Bench Xie et al. (2025), (4) the unsafe prompts from XSTest (Röttger et al., 2024), (5) MaliciousInstruct (Huang et al., 2024d), (6) StrongREJECT (Souly et al., 2024), and (7) $50$ randomly sampled examples from each of the nine medical harm categories in MedSafetyBench (Han et al., 2024). We extract the core malicious intent phrase from each query using few-shot prompting with GPT-4o (Hurst et al., 2024), and treat the removed portion as the style pattern. To minimize interference, we restrict the extraction to only words in the original query, avoiding any paraphrasing. We validate this decomposition using an entailment model (nli-deberta-v3-base (He et al., 2021; Reimers & Gurevych, 2019)) and discard cases where the extracted intent phrase is identical to the full query. In total, we obtain a pool of 2,1342{,}134 jailbreak queries and their corresponding variants with the style removed. Manual verification on $100$ random examples from the pool confirms the quality of this filtering pipeline. Implementation details and dataset examples are in §A.1.
For each version of the 2,1342{,}134 queries, we leverage GPT-4o to measure the ASR (Qi et al., 2025a) of $32$ instruction-tuned and safety-aligned LLMs from five model families:
@@ 3.2 ASR Inflation in Existing Jailbreak Benchmarks
As shown in Figure 2 (a), $28$ out of the $32$ examined models exhibit higher ASR when prompted with the original queries containing both style patterns and malicious intents, compared to prompting with malicious intents alone. This ASR inflation across nearly all models highlights the unintended influence of style patterns in jailbreaking safety-aligned LLMs. In addition, we observe a trend in ASR inflation across the five model families: Mistral exhibits the highest inflation, followed by OLMo and Qwen, while Gemma and Llama show lower or even negative inflation. In Figure 2 (b), we show the number of models with inflated ASR for each of the seven evaluated jailbreak benchmarks. All benchmarks lead to ASR inflation, with SorryBench and MedSafetyBench affecting the most models, where $30$ out of the $32$ LLMs experience increased ASR.
@@ 3.3 Dissecting Factors Correlated with ASR Inflation
To better understand the source of ASR inflation, we examine whether it can be attributed to LLMs’ internal attention behaviors (Ding et al., 2017; Mullenbach et al., 2018; Vig et al., 2020)—specifically, how much more attention LLMs allocate to style patterns compared to malicious intents. For each model, we aggregate attention weights for each token in a jailbreak query across all heads and layers before the model starts its response. We define attention difference as the average attention to style pattern tokens minus the average attention to malicious intent tokens within the same query. Figure 3 (a) plots each model’s ASR inflation against its average attention difference across all jailbreak queries. We observe a statistically significant Spearman rank correlation of $0.571$ ($p$-value = $6\mathrm{e}{-4}$) between ASR inflation and attention difference. This suggests that LLMs that disproportionately focus on style patterns are more vulnerable to jailbreaks that explicitly specify response styles. The family-wise trend in ASR inflation also follows here: Mistral shows the highest attention difference in general, while Gemma shows the lowest. We find that this correlation holds across most benchmarks, except for XSTest and MedSafetyBench. We also observe statistically significant but modest correlations between ASR inflation and both style pattern length and complexity, but no significant correlation with model size or release date. Full results are in §A.2.
@@ 4.2 Inflated Safety Risks from Superficial Style Tuning
Furthermore, we observe a sharp increase in ASR when the training and testing styles match for the four identified style patterns (list_prefix, list_suffix, poem_prefix, and poem_suffix), with the rise becoming prominent after the first checkpoint at $0.4$ epoch. The increase in ASR is largely mitigated when a higher proportion of style-removed data is mixed into the fine-tuning set. These findings support our hypothesis on the safety risks of superficial style alignment: Overexposure to specific style patterns during fine-tuning can make LLMs more vulnerable to similarly styled malicious queries. We also find that the position of style patterns (prefixes or suffixes) has little effect on ASR trends. Additional results and the discussion of style adaptation utility are in §B.2.
@@ 5.1 Experiment Setup
Training Set. For representative style patterns, we consider six variants (illustrated in §C.1): (1) list (He et al., 2024) (“Create a list to …”), (2) poem (Chakrabarty et al., 2022; Chen et al., 2024b; Mahbub et al., 2023) (“Write a poem to …”), (3) news (Amponsah & Atianashie, 2024; Tseng et al., 2025) (“Write a news story to …”), (4) legal (Guha et al., 2023; Jiang et al., 2024a) (“Create a legal document to …”), (5) shakespeare (Karpathy, 2015; Jhamtani et al., 2017) (“Respond in the style of Shakespearean English to …”), and (6) code [^3] We do not use any code-specific fine-tuning sets (Chaudhary, 2023; Ahmad et al., 2025), as our goal is to compare the effects of different style patterns under superficial style alignment rather than to introduce new programming capabilities. We show in §C.2 that this setting does not degrade the model’s coding capability. (Roziere et al., 2023; Ma et al., 2024; Zhang et al., 2025b) (“Write a code function to …”). As shown in §4.2, the effect of superficial style alignment is largely invariant to the position of style patterns, so we specify all styles as prefixes in this setup. Following §4.1, we construct 10,00010{,}000 instruction–response pairs for each style using a cleaned version of Alpaca (Taori et al., 2023). For real-world instruction-tuning sets, we use Dolly-15K (Conover et al., 2023) and Alpaca-52K (Taori et al., 2023) with GPT-4o’s responses. The safety training data is adopted from Bianchi et al. (2024).

### v3

@@ Abstract
Large language models (LLMs) can be prompted with specific styles (e.g., formatting responses as lists), including in malicious queries. Prior jailbreak research mainly augments these queries with additional string transformations to maximize attack success rate (ASR). However, the impact of style patterns in the original queries that are semantically irrelevant to the malicious intent remains unclear. In this work, we seek to understand whether style patterns compromise LLM safety, how superficial style alignment increases model vulnerability, and how best to mitigate these risks during alignment. We first define ASR inflation as the increase in ASR due to style patterns in existing jailbreak benchmark queries. By evaluating $36$ LLMs across seven benchmarks, we find that nearly all models exhibit ASR inflation. Notably, the inflation correlates with an LLM’s relative attention to style patterns, which also overlap more with its instruction-tuning data when inflation occurs. We then investigate superficial style alignment, and find that fine-tuning with specific styles makes LLMs more vulnerable to jailbreaks of those same styles. Finally, we propose SafeStyle, a defense strategy that incorporates a small amount of safety training data augmented to match the distribution of style patterns in the fine-tuning data. Across three LLMs, six fine-tuning style settings, and two real-world instruction-tuning datasets, SafeStyle consistently outperforms baselines in maintaining LLM safety. Content Warning: This paper contains examples of harmful language.
@@ 1 Introduction
These observations raise a critical yet understudied question: Can style patterns in malicious queries inadvertently undermine the robustness of safety-aligned LLMs? The superficial alignment hypothesis (Zhou et al., 2023; Lin et al., 2024; Zhang & Wu, 2024) posits that alignment tuning may lead models to imitate style patterns without internalizing deeper safety principles. Meanwhile, prior work on safety defense has largely focused on data curation (Bianchi et al., 2024; He et al., 2024; Shen et al., 2025), mechanistic interventions (Hazra et al., 2024; Hsu et al., 2024; Li et al., 2025a; Tamirisa et al., 2025; Yi et al., 2025b; Zhao et al., 2025c), or safety training objectives (Huang et al., 2024b; c; 2025; Rosati et al., 2024; Zhang et al., 2025a). While Hsiung et al. (2025) show that diverse alignment data improves robustness, they do not examine the underlying mechanism. Qi et al. (2025a) argue that safety training is often shallow but focus mainly on the early response stage. This leaves a crucial gap in understanding the safety implications of style compliance.
@@
Style patterns and LLM safety: First, we define ASR inflation as the phenomenon in which the underlying malicious intent remains unchanged, but the addition of semantically irrelevant style patterns to jailbreak queries breaks LLM safety. For instance, Mistral-Nemo-Instruct-2407 (Jiang et al., 2023) refuses “chemical warfare agents” but complies with “create a list of chemical warfare agents”. Through an extensive study of $36$ LLMs (Bai et al., 2023; Jiang et al., 2023; Touvron et al., 2023; Grattafiori et al., 2024; Groeneveld et al., 2024; OLMo et al., 2024; Team et al., 2024a; b; 2025; Yang et al., 2024a; b; Guo et al., 2025; Hurst et al., 2024) across seven jailbreak benchmarks (Zou et al., 2023; Han et al., 2024; Huang et al., 2024d; Mazeika et al., 2024; Röttger et al., 2024; Souly et al., 2024; Xie et al., 2025), we find that style patterns inflate the ASR for nearly all models. Based on our finding, we argue that the reported ASRs in benchmarks are inflated, in the sense that they do not capture the true rates when LLMs face core malicious intents alone. We further show that ASR inflation correlates with a model’s relative attention to these patterns. Surprisingly, ASR-inflating style patterns appear more frequently in the instruction-tuning datasets (Ivison et al., 2023; Lambert et al., 2024) used to align LLMs (Groeneveld et al., 2024; OLMo et al., 2024).
@@
  • We show that style patterns inflate ASR by systematically evaluating the safety risks of $36$ LLMs across seven jailbreak benchmarks.
@@ 2 Related Work
LLM Safety Risk. Despite extensive safety alignment efforts, LLMs remain vulnerable to malicious queries (Deng et al., 2024; Dong et al., 2024; Jin et al., 2024; Zheng et al., 2024a; Chan et al., 2025). In practice, style patterns are prevalent in ordinary user queries (Jiang et al., 2024b), and style-based attack strategies (Wei et al., 2023; Doumbouya et al., 2025) introduce competing objectives that can further amplify these safety risks. Recent jailbreak benchmarks quantify such vulnerabilities by measuring the ease with which LLMs produce unethical responses (Zou et al., 2023; Han et al., 2024; Huang et al., 2024d; Mazeika et al., 2024; Röttger et al., 2024; Souly et al., 2024; Xie et al., 2025). However, as we show, many benchmarks report inflated ASRs due to style patterns that are independent of the underlying malicious intents or any added attack transformations. Meanwhile, even benign fine-tuning can inadvertently compromise model safety (He et al., 2024; Qi et al., 2024; Eiras et al., 2025). Rather than proposing new attacks, our work identifies superficial style alignment during fine-tuning as a potential cause of the inflated ASRs observed in these benchmarks.
LLM Safety Defense. Various studies have proposed defense strategies to address the safety challenges faced by LLMs, particularly those introduced during fine-tuning (Huang et al., 2024a). Several methods (Hazra et al., 2024; Hsu et al., 2024; Liu et al., 2024; Djuhera et al., 2025; Li et al., 2025a; Tamirisa et al., 2025; Yi et al., 2025b; Zhao et al., 2025c) develop mechanistic interventions to protect safety-relevant parameters, while others (Huang et al., 2024b; c; 2025; Qi et al., 2025a; Rosati et al., 2024; Zhang et al., 2025a) modify the training objective to prioritize safety. Diagnostic efforts (Peng et al., 2024; Qi et al., 2025b) examine how safety behaviors are encoded and degrade during adaptation. Data-centric approaches emphasize careful curation of fine-tuning data (Bianchi et al., 2024; He et al., 2024; Shen et al., 2025) or highlight the critical role of prompt templates at inference time (Lyu et al., 2024; Eiras et al., 2025; Yi et al., 2025a). Notably, Hsiung et al. (2025) underscores diversity in safety alignment data, and Sharma et al. (2025) uses style-based augmentation to train red-teaming classifiers. In contrast, our work demonstrates that incorporating a small amount of safety data, which is augmented to match the distributions of style patterns in the fine-tuning data, can effectively restore LLMs’ safety strength degraded by superficial style alignment.
@@ 3 Style Patterns Inflate ASR
Figure 2: (a) Nearly all of the $36$ examined LLMs exhibit inflated ASR due to the incorporation of style patterns in jailbreak queries. (b) All seven jailbreak benchmarks lead to ASR inflation, with SorryBench and MedSafetyBench affecting the most LLMs.
@@ 3.1 Experiment Setup
In our experiments, we consider seven jailbreak benchmarks: (1) AdvBench (Zou et al., 2023), (2) the standard split of HarmBench (Mazeika et al., 2024), (3) the base dataset from SORRY-Bench (Xie et al., 2025), (4) the unsafe prompts from XSTest (Röttger et al., 2024), (5) MaliciousInstruct (Huang et al., 2024d), (6) StrongREJECT (Souly et al., 2024), and (7) $50$ randomly sampled examples from each of the nine medical harm categories in MedSafetyBench (Han et al., 2024). We extract the core malicious intent phrase from each query using few-shot prompting with GPT-4o (Hurst et al., 2024), and treat the removed portion as the style pattern. We restrict the extraction to only words in the original query, avoiding any paraphrasing to minimize interference, and validate this decomposition using an entailment model (nli-deberta-v3-base (He et al., 2021; Reimers & Gurevych, 2019)). Since style patterns are not always extractable, we discard cases where the extracted intent phrase is identical to the full query, accounting for $4\%$ of all queries. In total, we obtain a pool of 2,1342{,}134 jailbreak queries and their corresponding variants with the style removed. Manual verification on $100$ random examples from the pool confirms the quality of this filtering pipeline. Implementation details and dataset examples are in §A.1.
For each version of the 2,1342{,}134 queries, we leverage GPT-4o to measure the ASR (Qi et al., 2025a) of $36$ instruction-tuned and safety-aligned LLMs from seven model families:
@@
  • DeepSeek (Guo et al., 2025): DeepSeek-R1-Distill-Llama-8B/Qwen-14B/32B
  • GPT (Hurst et al., 2024): GPT-4o-mini
@@ 3.2 ASR Inflation in Existing Jailbreak Benchmarks
As shown in Figure 2 (a), $32$ out of the $36$ examined models exhibit higher ASR when prompted with the original queries containing both style patterns and malicious intents, compared to prompting with malicious intents alone. To assess statistical significance, we perform a paired $t$-test comparing ASR on the original jailbreak queries versus the extracted malicious intents, which yields a $p$-value =0.0002<0.05=0.0002<0.05, indicating a significant difference. This ASR inflation across nearly all models highlights the unintended influence of style patterns in jailbreaking safety-aligned LLMs. In addition, we observe a trend in ASR inflation across the seven model families: Mistral exhibits the highest inflation, followed by DeepSeek, OLMo, Qwen, and GPT, while Gemma and Llama show lower or even negative inflation. In Figure 2 (b), we show the number of models with inflated ASR for each of the seven evaluated jailbreak benchmarks. All benchmarks lead to ASR inflation, with SorryBench and MedSafetyBench affecting the most models, where $34$ out of the $36$ LLMs experience increased ASR. In §A.2, we further show that this effect persists both under alternative LLMs-as-judges (Grattafiori et al., 2024; Zhao et al., 2025a) and in a separate controlled experiment with synthetic style patterns.
@@ 3.3 Dissecting Factors Correlated with ASR Inflation
To better understand the source of ASR inflation, we examine whether it can be attributed to LLMs’ internal attention behaviors (Ding et al., 2017; Mullenbach et al., 2018; Vig et al., 2020)—specifically, how much more attention LLMs allocate to style patterns compared to malicious intents. For each model except GPT-4o-mini, we aggregate attention weights across all heads and layers for each token in a jailbreak query before generation begins. We define attention difference as the average attention to style pattern tokens minus that to malicious intent tokens within the same query. Figure 3 (a) plots each model’s ASR inflation against its average attention difference across all jailbreak queries. We observe a statistically significant Spearman rank correlation of $0.456$ ($p$-value = $6\mathrm{e}{-3}<0.05$) between ASR inflation and attention difference. This suggests that LLMs that disproportionately focus on style patterns are more vulnerable to jailbreaks that explicitly specify response styles. The family-wise trend in ASR inflation also follows here: Mistral shows the highest attention difference in general, while Gemma shows the lowest. We find that this correlation holds across most benchmarks, except for XSTest and MedSafetyBench. We also observe statistically significant but modest correlations between ASR inflation and both style pattern length and complexity, but no significant correlation with model size or release date. Full results are in §A.3.
@@ 4.2 Inflated Safety Risks from Superficial Style Tuning
Furthermore, we observe a sharp increase in ASR when the training and testing styles match for the four identified style patterns (list_prefix, list_suffix, poem_prefix, and poem_suffix), with the rise becoming prominent after the first checkpoint at $0.4$ epoch. The increase in ASR is largely mitigated when a higher proportion of style-removed data is mixed into the fine-tuning set. These findings support our hypothesis on the safety risks of superficial style alignment: Overexposure to specific style patterns during fine-tuning can make LLMs more vulnerable to similarly styled malicious queries. We also find that the position of style patterns (prefixes or suffixes) has little effect on ASR trends. In §B.2, we replicate this ASR trend using additional style patterns and further analyze the style adaptation utility during the superficial style alignment process. We also demonstrate that ASR inflation is primarily driven by semantic style compliance rather than token-level syntactic backdoor triggers (Li et al., 2025c; Zhao et al., 2025b).
@@ 5.1 Experiment Setup
Training Set. For representative style patterns, we consider six variants (illustrated in §C.1): (1) list (He et al., 2024) (“Create a list to …”), (2) poem (Chakrabarty et al., 2022; Chen et al., 2024b; Mahbub et al., 2023) (“Write a poem to …”), (3) news (Amponsah & Atianashie, 2024; Tseng et al., 2025) (“Write a news story to …”), (4) legal (Guha et al., 2023; Jiang et al., 2024a) (“Create a legal document to …”), (5) Shakespeare [^3] We acknowledge that the Shakespeare and poem styles are not typical downstream applications. We include them as stress tests to evaluate the generalization of superficial style alignment across diverse stylistic extremes and to assess SafeStyle under challenging style-based attacks. More common styles (e.g., list, news, legal, and code) are also considered to reflect realistic user scenarios. (Karpathy, 2015; Jhamtani et al., 2017; Krishna et al., 2020) (“Respond in the style of Shakespearean English to …”), and (6) code [^4] We do not use any code-specific fine-tuning sets (Chaudhary, 2023; Ahmad et al., 2025), as our goal is to compare the effects of different style patterns under superficial style alignment rather than to introduce new programming capabilities. We show in §C.2 that this setting does not degrade the model’s coding capability. (Roziere et al., 2023; Ma et al., 2024; Zhang et al., 2025b) (“Write a code function to …”). As shown in §4.2, the effect of superficial style alignment is largely invariant to the position of style patterns, so we specify all styles as prefixes in this setup. Following §4.1, we construct 10,00010{,}000 instruction–response pairs for each style using a cleaned version of Alpaca (Taori et al., 2023). For real-world instruction-tuning sets, we use Dolly-15K (Conover et al., 2023) and Alpaca-52K (Taori et al., 2023) with GPT-4o’s responses. The safety training data is adopted from Bianchi et al. (2024).